The animation in the simulation (linked below) is a simple depiction of the motion of an orange, which is treated as a particle. The simulation depicts two reference frames, which are labeled A and BRun the animation now. The motion is from the perspective of an observer (perhaps named Alex) at rest in frame A. Two things are being observed from frame A: the orange, and frame B. Here are some ideas to keep in mind about the simulation:


For the default values of vBA,x=4 m/s and vBA,y=0, what are the values of xPA (the x coordinate of the orange (the particle) as observed by Alex in frame A) at t=0

 and at t=8 s?




1
Expert's answer
2021-02-17T12:07:33-0500

The given parameters are:-

"v=4m\/s,y=0,x=0"

as we Know speed of light, "c=3\times 10^8m/s"


As there are two cordinate frames-


From the alex point of view

There was a transfomation in the cordinate of orange.


According to lorentz transformation,

"x'=\dfrac{x+vt}{\sqrt{1-\dfrac{v^2}{c^2}}}"


At time t=0, The cordinate of x is-


"=\dfrac{0+4\times 0}{\sqrt{1-\dfrac{v^2}{c^2}}}=0"


at time t=8s, The cordinate of x is:-

"=\dfrac{0+4\times 8}{\sqrt{1-\dfrac{(4)^2}{(3\times 10^8)^2}}}"


As "4<<3\times 10^8" So It can be neglected,

The cordinate of "x \text{ is } 32m."